OpenStudy (anonymous):

The sum of the interior angles of a pentagon is ______ degrees. a. 180 b. 360 c. 540 d. 720 The sum of the interior angles of an octagon is _______ degrees. a. 1080 b. 1440 c. 1800 d. 180

OpenStudy (anonymous):

sum of exterior angles=360 degree if no. of sides =n then sum of interior angles=180*n-360=180(n-2)